find the value(s) of t so that the distance from P(3,4) to R(t,8) is 4√2


1
Expert's answer
2022-03-03T17:32:45-0500

Let us find the values of "t" so that the distance from "P(3,4)" to "R(t,8)" is "4\sqrt{2}."

It follows that

"\sqrt{(t-3)^2+(8-4)^2}=4\sqrt{2},"

and hence

"(t-3)^2+16=32."


Therefore, "(t-3)^2=16," and thus "t-3=\pm 4."


We conclude that "t=-1" or "t=7."